bellapollan05:

Which number can each term of the equation be multiplied by to eliminate the fractions before solving? m – negative StartFraction 3 Over 4 EndFraction m minus StartFraction one-half EndFraction equals 2 StartFraction one-fourth EndFraction m. = 2 + m

jhonyy9:

\[m - \frac{ 3 }{ 4 }m - \frac{ 1 }{ 2 } = 2 \frac{ 1 }{ 4 }m,= 2 + m\] do you think it in this way ? but at the end with this virgule next the m and in front of equal sign so what wan be there ? please clarifie it